OpenStudy (anonymous):

what is the value of the expression 2.4 x 10^3? Please DO NOT give me the answer. I forgot how to work these out so explain the steps to me please!

OpenStudy (anonymous):

Please don't give me the answer D:

OpenStudy (ajprincess):

10^3=1000 so what do u get when u multiply 2.4 by 1000?

OpenStudy (anonymous):

Ohh so the next step is to multiply it?

OpenStudy (ajprincess):

ya 2.4*1000=?

OpenStudy (anonymous):

2,400? x-x Sorry, I'm rusty.

OpenStudy (ajprincess):

2.4=24/10 2.4*1000=(24/10)*1000=?

OpenStudy (anonymous):

I don't get it.. x.x

OpenStudy (ajprincess):